A circular coil with 100 turns and a radius of 0.10 m is placed in a uniform magnetic

field so that it is perpendicular to the plane of the coil. The field changes uniformly

from zero to 0.4T in 2 seconds. Calculate the induced emf.


1
Expert's answer
2022-06-15T07:03:43-0400

"\varepsilon=\frac{n \pi r^2B}t=0.63~V."